# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ F.3 Wind Turbine Power
Figure 11: Power curve of a Vestas V150-4.2 data in region 2 (blue bars) and the fitted polynomial of degree 6 (red function). More precisely, the interpolated power curve (red function), denoted here as $P^{\text{W}}_{R_{2}}$, determines the generated wind power $P^{\text{W}}$ depending on the wind speed $W$, in region 2.
To achieve a function that is a representation of the curve, the following form is used
Appendix G Details on Control Constraints
This provide the derivation of the control constraints given in Subsection 3.3 below.
Upper Bound
A first upper bound for values of the control process $A$ is obtained from a natural upper bound of the HTHX outlet temperature $T^{\text{HT,out}}_{n}$. This follows from relation (10) and its monotonicity with respect to the compressor shaft speed. Replacing $D_{n}$ by the maximum shaft speed $d_{\text{max}}$ of the HTHP and knowing that $T^{\text{HT,in}}_{n}=T^{\text{SG,out}}$ during the charging process, it holds $T^{\text{HT,out}}_{n}\leq\tau^{\text{HT,out}}_{\text{max}}:=F_{1}(T^{\text{SG,out}},\dot{m},T^{\text{LT,in}},d_{\text{max}})$ and relation (9) implies
A second upper bound follows from the fact that the charging factor $l^{\text{C}}$ is restricted to values in $[0,1]$. In charging operation, the average TES temperature increases, i.e. $R_{n}\leq R(t)$ for $t\in[t_{n},t_{n+1}]$. Since $T^{\text{HT,out}}_{n}$ and $T^{\text{SG,in}}$ are constants, (86) implies that the charging factor $l^{\text{C}}$ also increases. Thus, to ensure that $l^{\text{C}}\in[0,1]$ on $[t_{n},t_{n+1}]$, it is sufficient to require $l^{\text{C}}(t_{n+1})\leq 1$. This leads to the following upper bound for $A$ by inserting (3)-(8) into (86)
[…]
$\overline{a}(R_{n})=\min\{\overline{a}^{1},\overline{a}^{2}(R_{n})\}.$ (196)
Lower Bound
A first lower bound results from technical reasons, namely that the HTF temperature entering the heat pump must be bounded from above, by $\tau^{\text{HT,in}}_{\text{max}}$. During discharging it holds $T^{\text{HT,out}}_{n}=T^{\text{SG,in}}$ so that we obtain from relation (9)
A second lower bound follows, similar to the upper bound $\overline{a}^{2}$, from the fact that the discharging factor $l^{\text{D}}$ only takes values in $[0,1]$. During discharging, the TES temperature decreases with time, i.e. $R(t_{n+1})\leq R(t)$ for $t\in[t_{n},t_{n+1}]$. Since $T^{\text{HT,in}}_{n}$ and $T^{\text{SG,out}}$ are constants, (89) implies that the discharging factor $l^{\text{D}}$ increases. To ensure that $l^{\text{D}}\in[0,1]$ on $[t_{n},t_{n+1}]$, it is therefore sufficient to require $l^{\text{D}}(t_{n+1})\leq 1$. Inserting (3)-(8) into (89), the lower bound depending $R_{n}$ is obtained as follows:
Finally, the condition $R(t)\geq r_{\text{min}}=T^{\text{SG,out}}$ must fulfilled on $t\in[0,t_{\text{E}}]$. In discharging mode, $R$ is decreasing, so it is sufficient to require $R_{n+1}=R(t_{n+1})\geq T^{\text{SG,out}}$. Inserting $\underline{a}^{2}$ into (3) implies this condition. In summary, the following state-dependent lower bound of the control reads as
$\underline{a}(R_{n})=\max\{\underline{a}^{1},\underline{a}^{2}(R_{n})\}$ (199)
@@ H.1 Construction of Running Costs
During one time period $[t_{n},t_{n+1})$ the HTHP consumes a constant amount of energy $P^{\text{H}}_{n}$. For all times $t\in[t_{n},t_{n+1})$ the electricity consumption must be covered by the available wind energy $P^{\text{W}}(t)=P_{\text{WT}}(W(t))$ (which is a random variable) and the grid power $P^{\text{G}}(t)$, i.e.
@@
The expected operational cost at each time step $t\in[t_{n},t_{n+1})$ emerges from the price of the used grid power. Given state $x=(r,w,s)$ and action $a$ this yields the operational running costs $C_{n}(x,a)$ for the time interval $[t_{n},t_{n+1})$
$C_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\mathbb{E}_{n,x,a}[S(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{+}-\delta S_{sell}(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{-}]\mathrm{d}t$ (201)
with $\delta\in\{0,1\}$, $S_{sell}(t)=S(t)-\eta(t)$ and $\eta:[0,t_{\text{E}}]\to\mathbb{R}^{+}$. The amount of energy that needs to be bought is represented by the term $(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{+}$ and the amount energy being sold by $(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{-}$. Since $P^{\text{W}}(t)$ is bounded from above by $P^{\max}$, one can distinguish between two cases for the constant value of $P^{\text{H}}_{n}=\pi^{\text{H}}(T^{\text{HT,out}},T^{\text{HT,in}})$ as shown in Figure 12
Figure 12: Different cases of the constant HTHP energy consumption $P^{\text{H}}_{n}$. Case 1 ($P^{\text{H}}_{n}\geq P^{max}$): Due to the maximal rated power of the WT one always needs to buy additional electricity from the grid. Case 2 ($P^{\text{H}}_{n}<P^{max}$): One can find a wind speed $w^{*}\in R_{2}$ for which the electricity consumption $P^{\text{H}}_{n}$ of the HTHP is exactly covered. For all speeds below $w^{*}$ electricity from the grid needs to be bought. Above this value the WT produces an energy excess which can be sold. The right plot visualizes the buying and selling regions $R_{B}$ and $R_{S}$ in case two, respectively.
@@
If $P^{\text{H}}_{n}\geq P^{\max}$ we always need to buy additional electricity from the grid resulting in
@@
$C^{1}_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\mathbb{E}_{n,x,a}[S(t)(P^{\text{H}}_{n}-P^{\text{W}}(t))]\mathrm{d}t=\int_{t_{n}}^{t_{n+1}}P^{\text{H}}_{n}\mathbb{E}_{n,x,a}[S(t)]-\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))]\mathrm{d}t$ (203)
$=\int_{t_{n}}^{t_{n+1}}P^{\text{H}}_{n}m^{S}(\tau,y_{W},y_{S})-\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))]\mathrm{d}t.$ (204)
@@
For $P^{\text{H}}_{n}<P^{\max}$ there exists a $w^{*}\in R_{2}$ such that $P^{\text{W}}_{R_{2}}(w^{*})=P^{\text{H}}_{n}$, which separates the wind speeds into $R_{B}=R_{1}\cup R^{*}_{-}\cup R_{4}=(0,w_{in})\cup[w_{in},w^{*})\cup[w_{out},\infty)$, for which one has to buy from the grid and $R_{S}=R^{*}_{+}\cup R_{3}=[w^{*},w_{r})\cup[w_{r},w_{out})$, where one is able to sell unused energy. In our case the minimal electricity consumption of the HTHP is greater than the energy produced at cut-in of the WT we get $w^{*}\geq w_{in}$. Moreover, since $P^{\text{H}}_{n}<P^{\max}$ we also have $w^{*}<w_{r}$ and $w^{*}\in R_{2}=[w_{in},w_{r})$. Incorporating this to the running costs leads to
$C^{2}_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)(P^{\text{H}}_{n}-P^{\text{W}}(t))+\mathbbm{1}_{R_{S}}(W(t))\delta S_{sell}(t)(P^{\text{H}}_{n}-P^{\text{W}}(t))]\mathrm{d}t$ (205)
Note however that $w^{*}$ is a zero of $P^{\text{W}}_{R_{2}}(w)-P^{\text{H}}_{n}$ which is a polynomial of degree 6 and has at most 6 zeros. By construction there will be a zero in region 2, which need to find. Since no closed-form solution for $w^{*}$ exists, it is computed numerically.
@@
A detailed look at the cost functions (204) and (LABEL:Cost_2) reveals that one needs to find the expected values for
$\begin{split}&\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)P^{\text{W}}(t))],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)P^{\text{W}}(t))],\\ &\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))P^{\text{W}}(t)]\end{split}$ (207)
@@
Observing that $R_{B}\cap R_{2}=[w_{in},w^{*})=R^{*}_{-},~{}R_{S}\cap R_{2}=[w^{*},w_{r})=R^{*}_{+},~{}R_{B}\cap R_{3}=\emptyset,~{}R_{S}\cap R_{3}=R_{3}$, $\mathbbm{1}_{R_{B}}=\mathbbm{1}_{R_{1}}+\mathbbm{1}_{R_{-}^{*}}+\mathbbm{1}_{R_{4}}$ and $\mathbbm{1}_{R_{S}}=\mathbbm{1}_{R_{+}^{*}}+\mathbbm{1}_{R_{3}}$ yields for the expressions in (207)
$\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))]$ $=\sum_{k=0}^{6}a_{k}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{2}}(W(t))W(t)^{k}S(t)]+P^{\max}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{3}}(W(t))S(t)],$ (209)
$\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)P^{\text{W}}(t))]$ $=\sum_{k=0}^{6}a_{k}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R^{*}_{-}}(W(t))W(t)^{k}S(t)],$ (210)
$\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)P^{\text{W}}(t))]$ $=\sum_{k=0}^{6}a_{k}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R^{*}_{+}}(W(t))W(t)^{k}S(t)]+P^{\max}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{3}}(W(t))S(t)],$ (211)
$\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)]$ $=\mathbb{E}_{n,x,a}[(\mathbbm{1}_{R_{1}}(W(t))+\mathbbm{1}_{R^{*}_{-}}(W(t))+\mathbbm{1}_{R_{4}}(W(t)))S(t)],$ (212)
$\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)]$ $=\mathbb{E}_{n,x,a}[\mathbbm{1}_{R^{*}_{+}}(W(t))S(t)]+\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{3}}(W(t))S(t)],$ (213)
$\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))P^{\text{W}}(t)]$ $=\sum_{k=0}^{6}a_{k}\mathbb{E}_{w}[\mathbbm{1}_{R^{*}_{+}}(W(t))W(t)^{k}]+P^{\max}\mathbb{E}_{w}[\mathbbm{1}_{R_{3}}(W(t))].$ (214)
@@
$E(t,k,R)=\mathbb{E}_{n,x,a}[\mathbbm{1}_{R}(W(t))W(t)^{k}S(t)]\quad\text{ and }\quad E_{0}(t,k,R)=\mathbb{E}_{w}[\mathbbm{1}_{R}(W(t))W(t)^{k}]$ (215)
with $R\in\{R_{2},R_{3},R_{4},R^{*}_{-},R_{B,S},R^{*}_{+},R_{S}\}$ and $k\in\mathbb{N}$. Furthermore, $E_{0}(t,k,R)$ is the special case of $E(t,k,R)$, when the price is replaced with $S(t)=1$. Hence, it is sufficient to find a closed-form expression for $E(t,k,R)$ in order to get the corresponding expected values in $C_{n}^{1}$ and $C_{n}^{2}$, respectively. In total we get that cost function $C_{n}$ is given by
$C_{n}(x,a)=\begin{cases}&C^{1}_{n}(x,a),\quad P^{\text{H}}_{n}\geq P^{\max},\\ &C^{2}_{n}(x,a),\quad P^{\text{H}}_{n}<P^{\max},\end{cases}$
where $C_{n}^{1}$ and $C_{n}^{2}$ given in (216),(217). They can be rewritten as
$C_{n}^{1}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\Psi^{1}(t,x,a)\mathrm{d}t$ (216)
$C^{2}_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\Psi^{2}(t,x,a)\mathrm{d}t$ (217)
with functions
$\Psi^{1}(t,x,a)=$ $P^{\text{H}}_{n}m^{S}(\tau,y_{W},y_{S})-\sum_{k=0}^{6}a_{k}E(t,k,R_{2})-P^{\max}E(t,0,R_{3})$ (218)
$\Psi^{2}(t,x,a)=$ $P^{\text{H}}_{n}(E(t,0,R_{B})+\delta[E(t,0,R_{S})-\eta(t)E_{0}(t,0,R_{S})])$ (219)
$-\sum_{k=0}^{6}a_{k}\big{(}E(t,k,R^{*}_{-})+\delta[E(t,k,R^{*}_{+})-\eta(t)E_{0}(t,k,R^{*}_{+})]\big{)}$ (220)
$-\delta P^{\max}[E(t,0,R_{3})-\eta(t)E_{0}(t,0,R_{3})].$ (221)
To keep expressions compact note that for regions $R_{B}$ and $R_{S}$ we have
$E(t,0,R_{B})=E(t,0,R_{1})+E(t,0,R^{*}_{-})+E(t,0,R_{4}),$ (222)
$E(t,0,R_{S})=E(t,0,R^{*}_{+})+E(t,0,R_{3}),$ (223)
$E_{0}(t,0,R_{S})=E_{0}(t,0,R^{*}_{+})+E_{0}(t,0,R_{3}).$ (224)
@@
The calculation of $E(t,k,R)$ with $R=[a,b),a<b\in\mathbb{R}$ requires the joint density $f_{WS}$ of $W(t)$ and $S(t)$. For the sake of simplicity in the calculation we write $m^{W},m^{S}$ for $m^{W}(\tau,y_{W}),m^{S}(\tau,y_{W},y_{S})$ and $\Sigma_{W}^{2},\Sigma_{S}^{2}$ for $\Sigma_{W}^{2}(\tau),\Sigma_{S}^{2}(\tau)$ leading to
$E(t,k,R)=\int_{(0,\infty)}\int_{\mathbb{R}}\mathbbm{1}_{R}(w)w^{k}s\,f_{WS}(w,s)\mathrm{d}s\,\mathrm{d}w=\int_{a}^{b}\int_{\mathbb{R}}w^{k}s\,f_{WS}(w,s)\mathrm{d}s\,\mathrm{d}w$ (225)
$=\int_{a}^{b}\int_{\mathbb{R}}\,\tfrac{w^{k}s}{2\pi\Sigma_{W},\Sigma_{S}\sqrt{1-\rho^{2}}w}\exp\bigg{\{}-\tfrac{1}{2(1-\rho^{2})}\bigg{[}\tfrac{(\log(w)-m^{W})^{2}}{\Sigma_{W}^{2}}-2\rho\tfrac{(\log(w)-m^{W})(s-m^{S})}{\Sigma_{W}\Sigma_{S}}+\tfrac{(s-m^{S})^{2}}{\Sigma_{S}^{2}}\bigg{]}\bigg{\}}\mathrm{d}s\,\mathrm{d}w.$ (226)
Using the following substitutions $u=u(w)=\frac{\log(w)-m^{W}}{\Sigma_{W}},v=v(s)=\frac{s-m^{S}}{\Sigma_{S}}$ with
@@
$E(t,k,R)$ $=\int_{u(a)}^{u(b)}\int_{\mathbb{R}}\tfrac{w^{k}s}{2\pi\sqrt{1-\rho^{2}}}\exp\big{\{}-\tfrac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\big{\}}\mathrm{d}v\,\mathrm{d}u$ (228)
$=\int_{u(a)}^{u(b)}e^{k(\Sigma_{W}u+m^{W})}\int_{\mathbb{R}}\tfrac{\Sigma_{S}v+m^{S}}{2\pi\sqrt{1-\rho^{2}}}\exp\big{\{}-\tfrac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\big{\}}\mathrm{d}v\,\mathrm{d}u$ (229)
$=\int_{u(a)}^{u(b)}e^{k(\Sigma_{W}u+m^{W})}\int_{\mathbb{R}}\tfrac{\Sigma_{S}v+m^{S}}{2\pi\sqrt{1-\rho^{2}}}\exp\big{\{}-\tfrac{u^{2}-\rho^{2}u^{2}+\rho^{2}u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\big{\}}\mathrm{d}v\,\mathrm{d}u$ (230)

### v3

@@ G.3 Wind Turbine Power
Figure 11: Power curve data of a Vestas V150-4.2 in region 2 (blue bars) and the fitted polynomial of degree 6 (red). More precisely, the interpolated power curve (red), denoted here as $P^{\text{W}}_{R_{2}}$, determines the generated wind power $P^{\text{W}}$ depending on the wind speed $W$, in region 2.
@@ H.1 Construction of Running Costs
During one time period $t_{n},t_{n+1})$ the HTHP consumes a constant amount of energy $P^{\text{H}}_{n}$. For all times $t\in[t_{n},t_{n+1})$ the electricity consumption must be covered by the available wind energy $P^{\text{W}}(t)=P_{\text{WT}}(W(t))$ (which is a random variable) and the grid power $P^{\text{G}}(t)$, i.e.,
@@
The expected operational cost at each time step $t\in[t_{n},t_{n+1})$ emerges from the price of the used grid power. Given state $x=(r,w,s)$ and action $a$ this yields the operational running costs $C_{n}(x,a)$ for the time interval $t_{n},t_{n+1})$
$C_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\mathbb{E}_{n,x,a}[S(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{+}-\zeta S_{sell}(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{-}]\mathrm{d}t$ (161)
with $\zeta\in\{0,1\}$, $S_{sell}(t)=S(t)-\eta(t)$ and $\eta:[0,t_{\text{E}}]\to\mathbb{R}^{+}$. The amount of energy that needs to be bought is represented by the term $(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{+}$ and the amount energy being sold by $(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{-}$. Since $P^{\text{W}}(t)$ is bounded from above by $P^{\max}$, two cases are distinguished for the constant value of $P^{\text{H}}_{n}=\pi^{\text{H}}(T^{\text{HT,out}},T^{\text{HT,in}})$ as shown in Figure 12.
Figure 12: Different cases of the constant HTHP energy consumption $P^{\text{H}}_{n}$. Case 1 ($P^{\text{H}}_{n}\geq P^{max}$): Due to the maximal rated power of the WT one always needs to buy additional electricity from the grid. Case 2 ($P^{\text{H}}_{n}<P^{max}$): One can find a wind speed $w^{*}\in R_{2}$ for which the electricity consumption $P^{\text{H}}_{n}$ of the HTHP is exactly covered. For all speeds below $w^{*}$ electricity from the grid needs to be bought. Above this value the WT produces an energy excess, which can be sold. The right plot visualizes the buying and selling regions $R_{B}$ and $R_{S}$, respectively.
@@
If $P^{\text{H}}_{n}\geq P^{\max}$ we always need to buy additional electricity from the grid, resulting in
@@
$C^{1}_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\mathbb{E}_{n,x,a}[S(t)(P^{\text{H}}_{n}-P^{\text{W}}(t))]\mathrm{d}t=\int_{t_{n}}^{t_{n+1}}P^{\text{H}}_{n}\mathbb{E}_{n,x,a}[S(t)]-\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))]\mathrm{d}t$
$=\int_{t_{n}}^{t_{n+1}}P^{\text{H}}_{n}m^{S}(\tau,y_{W},y_{S})-\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))]\mathrm{d}t,$ (163)
@@
For $P^{\text{H}}_{n}<P^{\max}$ there exists a $w^{*}\in R_{2}$ such that $P^{\text{W}}_{R_{2}}(w^{*})=P^{\text{H}}_{n}$, which separates the wind speeds into $R_{B}=R_{1}\cup R^{*}_{-}\cup R_{4}=(0,w_{\text{in}})\cup[w_{\text{in}},w^{*})\cup[w_{\text{out}},\infty)$, meaning to purchase power from the grid and $R_{S}=R^{*}_{+}\cup R_{3}=[w^{*},w_{r})\cup[w_{r},w_{\text{out}})$, where it is able to sell unused electricity. In our case, the minimal electricity consumption of the HTHP is greater than the power generated at cut-in of the WT, we get $w^{*}\geq w_{\text{in}}$. Moreover, since $P^{\text{H}}_{n}<P^{\max}$, we also have $w^{*}<w_{r}$ and $w^{*}\in R_{2}=[w_{\text{in}},w_{r})$. Incorporating this to the running costs leads to
$\begin{split}C^{2}_{n}(x,a)&=\int_{t_{n}}^{t_{n+1}}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)(P^{\text{H}}_{n}-P^{\text{W}}(t))\\ &\hskip 8.5359pt+\mathbbm{1}_{R_{S}}(W(t))\zeta S_{sell}(t)(P^{\text{H}}_{n}-P^{\text{W}}(t))]\mathrm{d}t.\end{split}$ (164)
However, $w^{*}$ is a zero of $P^{\text{W}}_{R_{2}}(w)-P^{\text{H}}_{n}$, which is a polynomial of degree 6 and has at most 6 zeros. By construction, there is a zero in region 2 that needs to be found. Since no closed-form solution for $w^{*}$ exists, it is computed numerically.
@@
A detailed look at the cost functions (163) and (164) shows that it is required to determine the expected values for
$\begin{split}&\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)P^{\text{W}}(t))],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)P^{\text{W}}(t))],\\ &\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)],\quad\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))P^{\text{W}}(t)],\end{split}$ (165)
@@
Observing that $R_{B}\cap R_{2}=[w_{\text{in}},w^{*})=R^{*}_{-},~R_{S}\cap R_{2}=[w^{*},w_{r})=R^{*}_{+},~R_{B}\cap R_{3}=\emptyset,~R_{S}\cap R_{3}=R_{3}$, $\mathbbm{1}_{R_{B}}=\mathbbm{1}_{R_{1}}+\mathbbm{1}_{R_{-}^{*}}+\mathbbm{1}_{R_{4}}$ and $\mathbbm{1}_{R_{S}}=\mathbbm{1}_{R_{+}^{*}}+\mathbbm{1}_{R_{3}}$ yields for the expressions in (165)
$\begin{split}\mathbb{E}_{n,x,a}[S(t)P^{\text{W}}(t))]&=\sum_{k=0}^{6}a_{k}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{2}}(W(t))W(t)^{k}S(t)]\\ &\hskip 8.5359pt+P^{\max}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{3}}(W(t))S(t)],\\ \mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)P^{\text{W}}(t))]&=\sum_{k=0}^{6}a_{k}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R^{*}_{-}}(W(t))W(t)^{k}S(t)],\\ \mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)P^{\text{W}}(t))]&=\sum_{k=0}^{6}a_{k}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R^{*}_{+}}(W(t))W(t)^{k}S(t)]\\ &\hskip 8.5359pt+P^{\max}\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{3}}(W(t))S(t)],\\ \mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{B}}(W(t))S(t)]&=\mathbb{E}_{n,x,a}[(\mathbbm{1}_{R_{1}}(W(t))+\mathbbm{1}_{R^{*}_{-}}(W(t))+\mathbbm{1}_{R_{4}}(W(t)))S(t)],\\ \mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))S(t)]&=\mathbb{E}_{n,x,a}[\mathbbm{1}_{R^{*}_{+}}(W(t))S(t)]+\mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{3}}(W(t))S(t)],\\ \mathbb{E}_{n,x,a}[\mathbbm{1}_{R_{S}}(W(t))P^{\text{W}}(t)]&=\sum_{k=0}^{6}a_{k}\mathbb{E}_{w}[\mathbbm{1}_{R^{*}_{+}}(W(t))W(t)^{k}]+P^{\max}\mathbb{E}_{w}[\mathbbm{1}_{R_{3}}(W(t))].\end{split}$ (167)
@@
$E(t,k,R)=\mathbb{E}_{n,x,a}[\mathbbm{1}_{R}(W(t))W(t)^{k}S(t)]\quad\text{ and }\quad E_{0}(t,k,R)=\mathbb{E}_{w}[\mathbbm{1}_{R}(W(t))W(t)^{k}],$ (168)
with $R\in\{R_{2},R_{3},R_{4},R^{*}_{-},R_{B,S},R^{*}_{+},R_{S}\}$ and $k\in\mathbb{N}$. Furthermore, $E_{0}(t,k,R)$ is the special case of $E(t,k,R)$, when the price is replaced with $S(t)=1$. Hence, it is sufficient to find a closed-form expression for $E(t,k,R)$ in order to get the corresponding expected values in $C_{n}^{1}$ and $C_{n}^{2}$, respectively. In total, the cost function $C_{n}$ is given by
$C_{n}(x,a)=\begin{cases}&C^{1}_{n}(x,a),\quad P^{\text{H}}_{n}\geq P^{\max},\\ &C^{2}_{n}(x,a),\quad P^{\text{H}}_{n}<P^{\max},\end{cases}$ (169)
with $C_{n}^{1}$ and $C_{n}^{2}$ specified as
$C_{n}^{1}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\Psi^{1}(t,x,a)\mathrm{d}t,$ (170)
$C^{2}_{n}(x,a)$ $=\int_{t_{n}}^{t_{n+1}}\Psi^{2}(t,x,a)\mathrm{d}t,$ (171)
with the functions
$\Psi^{1}(t,x,a)=$ $P^{\text{H}}_{n}m^{S}(\tau,y_{W},y_{S})-\sum_{k=0}^{6}a_{k}E(t,k,R_{2})-P^{\max}E(t,0,R_{3}),$ (172)
$\Psi^{2}(t,x,a)=$ $P^{\text{H}}_{n}(E(t,0,R_{B})+\zeta[E(t,0,R_{S})-\eta(t)E_{0}(t,0,R_{S})])$
$-\sum_{k=0}^{6}a_{k}\big(E(t,k,R^{*}_{-})+\zeta[E(t,k,R^{*}_{+})-\eta(t)E_{0}(t,k,R^{*}_{+})]\big)$ (173)
$-\zeta P^{\max}[E(t,0,R_{3})-\eta(t)E_{0}(t,0,R_{3})].$
To keep expressions compact, note that for regions $R_{B}$ and $R_{S}$, we have
$\begin{split}&E(t,0,R_{B})=E(t,0,R_{1})+E(t,0,R^{*}_{-})+E(t,0,R_{4}),\\ &E(t,0,R_{S})=E(t,0,R^{*}_{+})+E(t,0,R_{3}),\\ &E_{0}(t,0,R_{S})=E_{0}(t,0,R^{*}_{+})+E_{0}(t,0,R_{3}).\end{split}$ (174)
@@
The calculation of $E(t,k,R)$ with $R=[a,b),a<b\in\mathbb{R}$ requires the joint density $f_{WS}$ of $W(t)$ and $S(t)$. For the sake of simplicity, we write $m^{W},m^{S}$ for $m^{W}(\tau,y_{W}),m^{S}(\tau,y_{W},y_{S})$ and $\Sigma_{W}^{2},\Sigma_{S}^{2}$ for $\Sigma_{W}^{2}(\tau),\Sigma_{S}^{2}(\tau)$ leading to
$\begin{split}E(t,k,R)&=\int_{(0,\infty)}\int_{\mathbb{R}}\mathbbm{1}_{R}(w)w^{k}s\,f_{WS}(w,s)\mathrm{d}s\,\mathrm{d}w=\int_{a}^{b}\int_{\mathbb{R}}w^{k}s\,f_{WS}(w,s)\mathrm{d}s\,\mathrm{d}w\\ &=\int_{a}^{b}\int_{\mathbb{R}}\,\tfrac{w^{k}s}{2\pi\Sigma_{W},\Sigma_{S}\sqrt{1-\rho^{2}}w}\exp\bigg\{-\tfrac{1}{2(1-\rho^{2})}\bigg[\tfrac{(\log(w)-m^{W})^{2}}{\Sigma_{W}^{2}}\\ &\hskip 8.5359pt-2\rho\tfrac{(\log(w)-m^{W})(s-m^{S})}{\Sigma_{W}\Sigma_{S}}+\tfrac{(s-m^{S})^{2}}{\Sigma_{S}^{2}}\bigg]\bigg\}\mathrm{d}s\,\mathrm{d}w.\end{split}$ (175)
Using the substitutions $u=u(w)=\frac{\log(w)-m^{W}}{\Sigma_{W}},v=v(s)=\frac{s-m^{S}}{\Sigma_{S}}$ with
@@
$E(t,k,R)$ $=\int_{u(a)}^{u(b)}\int_{\mathbb{R}}\tfrac{w^{k}s}{2\pi\sqrt{1-\rho^{2}}}\exp\big\{-\tfrac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\big\}\mathrm{d}v\,\mathrm{d}u$
$=\int_{u(a)}^{u(b)}e^{k(\Sigma_{W}u+m^{W})}\int_{\mathbb{R}}\tfrac{\Sigma_{S}v+m^{S}}{2\pi\sqrt{1-\rho^{2}}}\exp\big\{-\tfrac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\big\}\mathrm{d}v\,\mathrm{d}u$
$=\int_{u(a)}^{u(b)}e^{k(\Sigma_{W}u+m^{W})}\int_{\mathbb{R}}\tfrac{\Sigma_{S}v+m^{S}}{2\pi\sqrt{1-\rho^{2}}}\exp\big\{-\tfrac{u^{2}-\rho^{2}u^{2}+\rho^{2}u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\big\}\mathrm{d}v\,\mathrm{d}u$